For the phosphite ion, (PO3)^3- :
(a) giving all the steps, write the Lewis structure.
(b) calculate the formal charges on O atoms and P atom.
(c) draw the shape of the ion.
1
Expert's answer
2020-08-27T05:05:09-0400

Solution.

Since phosphorus is in the 5 (15) group of the periodic table, it has 5 electrons on the outer electron layer. Oxygen is in the 6(16) group of the periodic table, so it has 6 electrons on the outer layer. Since there are three oxygen atoms in this ion, the total will be 18 electrons from the oxygen atoms. Also, this ion has a charge. Since it is negative, we must add 3 more electrons (the value of the ion charge).

"5 + (6 \times 3) + 3 = 26 \ e"



If we look at the structural formula of an ion, then the oxygen atom is surrounded by 7 electrons, and the number of electrons in the free oxygen atom is 6, so the formal charge of oxygen is -1. And so it is with the others. Phosphorus is surrounded by 3 electrons that bind it to oxygen atoms and an unshielded electron pair, and the number of electrons in a free phosphorus atom is 5.Hence, the formal charge of phosphorus is 0.

Answer:

Since phosphorus is in the 5 (15) group of the periodic table, it has 5 electrons on the outer electron layer. Oxygen is in the 6(16) group of the periodic table, so it has 6 electrons on the outer layer. Since there are three oxygen atoms in this ion, the total will be 18 electrons from the oxygen atoms. Also, this ion has a charge. Since it is negative, we must add 3 more electrons (the value of the ion charge).

"5 + (6 \times 3) + 3 = 26 \ e"



If we look at the structural formula of an ion, then the oxygen atom is surrounded by 7 electrons, and the number of electrons in the free oxygen atom is 6, so the formal charge of oxygen is -1. And so it is with the others. Phosphorus is surrounded by 3 electrons that bind it to oxygen atoms and an unshielded electron pair, and the number of electrons in a free phosphorus atom is 5.Hence, the formal charge of phosphorus is 0.